### v3

# EFRame: Deeper Reasoning via Exploration-Filter-Replay Reinforcement Learning Framework

Abstract
Recent advances in reinforcement learning (RL) have significantly enhanced the reasoning capabilities of large language models (LLMs). Group Relative Policy Optimization (GRPO), an efficient variant of PPO that lowers RL’s computational cost, still faces limited exploration, low sample efficiency and instability, constraining its performance on complex reasoning tasks. To address these limitations, we introduce EFRame, an Exploration-Filter-Replay framework that systematically augments GRPO along three critical dimensions. EFRame performs additional rollouts to explore high-quality trajectories, applies online filtering to eliminate low-quality samples that introduce noise and variance, and leverages experience replay to repeatedly exploit rare but informative samples. EFRame establishes a complete and stable learning cycle, guiding the model through a structured transition from exploration to convergence. Our experiments across a variety of reasoning benchmarks demonstrate that EFRame not only improves the robustness and efficiency of training, but also enables access to deeper reasoning capabilities that remain unattainable under vanilla GRPO. Furthermore, EFRame not only enables fine-grained categorization of training samples for deeper insight into their contributions, but also introduces an efficient and precise mechanism for entropy control, which is critical for balancing exploration and convergence in RL training. Our code is available at https://github.com/597358816/EFRame.
1 Introduction
Reinforcement learning (RL) has played a pivotal role in the post-training of large language models (LLMs) GLM et al. (2024); Touvron et al. (2023). In the early stages, approaches such as Proximal Policy Optimization (PPO) and Direct Preference Optimization (DPO) were widely adopted under the Reinforcement Learning from Human Feedback (RLHF), effectively aligning model outputs with human preferences through preference-based rewards Schulman et al. (2017); Rafailov et al. (2023); Zhong et al. (2024); Wang et al. (2024b). More recently, the focus has shifted towards Reinforcement Learning with Verifiable Rewards (RLVR), which seeks to provide more consistent and scalable supervision by leveraging structured, verifiable reward signals Mroueh (2025). Within this paradigm, DeepSeek-R1 introduced Group Relative Policy Optimization (GRPO), a lightweight and efficient variant of PPO that computes relative rewards within rollout groups to eliminate the need for the critic model Shao et al. (2024); Liu et al. (2024); Guo et al. (2025). GRPO has demonstrated remarkable improvements in training efficiency and has shown promising potential in enhancing the reasoning capabilities of LLMs.
RLVR relies on extracting reward signals from the model’s own pass@$k$ outputs to improve its pass@1 performance. However, this design inherently couples final performance to the choice of $k$, as the reward signal vanishes when all sampled responses are either correct or incorrect, making it impossible for the model to extract useful knowledge from such data. This constrains RLVR’s learning capacity to the base model’s pass@$k$, limiting further exploration and improvement Yue et al. (2025). Although increasing the sampling number $k$ per prompt may seem like a straightforward solution, it typically yields only marginal improvements while significantly reducing sample efficiency Lin et al. (2025). This results in GRPO exhibiting limited exploration, and low sample efficiency, and indirectly leads to training instability in the later stages due to the disappearance of reward signals Yu et al. (2025); Li et al. (2025).
To address the aforementioned issues, we propose EFRame: an Exploration-Filter-Replay RL Framework. For challenging prompts that cannot be effectively handled by regular rollouts, EFRame introduces an additional rollout phase to enable targeted exploration. During this process, the model generates a small number of high-quality positive samples along with a large number of low-quality negative samples. To reduce training variance and improve efficiency, we apply an online filtering mechanism to discard low-quality samples. Meanwhile, high-advantage samples are stored in a replay buffer and reused during training to amplify their impact and enhance model stability.
The contributions of EFRame are threefold:
  1. 1.
    Enhanced High-Difficulty Reasoning: EFRame significantly improves the performance of GRPO on challenging reasoning tasks. By integrating exploration, filtering, and replay into a unified framework, it enables more complete and deeper reasoning processes in large language models.
  2. 2.
    Sample-Type-Aware Training: EFRame introduces a principled categorization of training samples to better understand their distinct roles in learning. It demonstrates that high-advantage, low-probability positive samples are crucial for triggering exploration and unlocking deeper knowledge, whereas low-advantage negative samples tend to drive premature convergence to suboptimal solutions and should be suppressed.
  3. 3.
    Precise and Efficient Entropy Control: EFRame offers a practical and effective approach to entropy control by jointly tuning the additional rollout temperature and the number of replayed samples. This mechanism provides more flexible and stable entropy regulation than traditional reward-based entropy regularization, supporting both efficient exploration and stable convergence.
EFRame provides a systematic solution to the inherent trade-offs among training stability, efficiency, and exploration capacity in reinforcement learning. By enabling the model to effectively identify and leverage deeper latent knowledge, EFRame enhances the expressiveness and learning potential of large language models. Through the coordinated integration of targeted exploration, online sample filtering, and experience replay, the framework not only addresses the limitations of GRPO in handling complex reasoning tasks but also establishes a more principled and controllable training paradigm. Empirical evaluations confirm the efficacy of EFRame: on the challenging Geometry3K benchmark, it outperforms GRPO by 36.5%, and it achieves consistent gains across a wide range of multimodal and mathematical reasoning benchmarks, demonstrating strong generalizability and robustness.
2 Related Work
The early application of RL in LLMs and MLLMs OpenAI (2023); Team et al. (2024); Zhu et al. (2023); Wei et al. (2023); Liu et al. (2023) was primarily reflected in RLHF, where algorithms such as PPO were employed to align language model outputs with human preferences and Direct Preference Optimization (DPO) Rafailov et al. (2024) bypass the need for reward modeling and instead optimize the model policy directly from preference data. Recently, RL has advanced toward integrating rule-based reward mechanisms with sophisticated optimization techniques, as exemplified by models such as DeepSeek-R1 Guo et al. (2025) and Kimi-1.5 Team et al. (2025). As a lightweight variant of PPO, GRPO Shao et al. (2024) significantly improves training efficiency, but at the cost of increased instability. A series of subsequent studies have introduced optimizations from GRPO. These include gradient loss refinements, as demonstrated in DAPO Yu et al. (2025) and Dr.GRPO Liu et al. (2025); online filtering methods like Meng et al. (2025); advantage enhancement techniques, exemplified by VL-Rethinker Wang et al. (2025a) and Luffy Yan et al. (2025); clip acceleration methods such as CPPO Lin et al. (2025); curriculum learning strategies explored in Chen et al. (2025b; c); and unsupervised RL such as Zhang et al. (2025); Wei et al. (2025a); Zhao et al. (2025); Agarwal et al. (2025). While existing methods typically focus on individual components, we systematically review the fundamental limitations of GRPO and propose a comprehensive framework with an integrated solution that addresses three critical aspects: exploration capacity, training stability, and sample efficiency. This helps create a more robust and effective GRPO-based training paradigm.
3 Overview of GRPO
3.1 Gradient Loss
GRPO has been employed to enhance the reasoning capabilities of large language models, particularly in mathematical problem-solving. Compared to PPO, GRPO improves computational efficiency by eliminating the need for a separate critic model. Instead, it estimates the advantage directly using group-normalized rewards. Given a question $q$ and a corresponding verifiable answer $a$, GRPO generates a set of responses $O=\{o_{i}\}_{i=1}^{G}$ using the policy model $\pi_{\theta_{old}}$, then optimizes the policy model by maximizing the following objective:
$\begin{split}\mathcal{J}_{GRPO}(\theta)&=\mathbb{E}_{(q,a)\sim\mathcal{D},\ O=\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}(\cdot|q)}}\\ &\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\left\{\min\left[r_{i,t}(\theta)\hat{A}_{i,t},\ \operatorname{clip}(r_{i,t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i,t}\right]-\beta D_{KL}[\pi_{\theta}||\pi_{ref}]\right\},\end{split}$ (1)
where $r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i}<t)}{\pi_{\theta_{old}}(o_{i,t}|q,o_{i}<t)}$ and $\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G})}$.
3.2 Limitations of GRPO
Although GRPO has demonstrated strong performance in various settings, it encounters several practical challenges when applied to complex reasoning tasks. In the following discussion, we highlight three main issues: limited exploration capacity, training collapse, and low sample efficiency. These limitations reveal the need for more robust and adaptive RL strategies.
Limited Exploration Capacity. GRPO lacks effective mechanisms to promote exploration, leading to persistently low entropy levels in the early stages of training and an increased risk of convergence to local optima. Moreover, it does not differentiate between prompts of varying difficulty, which limits its ability to stimulate deeper reasoning capabilities. As a consequence, only superficial knowledge is activated, while deeper knowledge remains underutilized.
((a))
((b))
((c))
Figure 1: Increasing the sampling number to $G=20$ provides a performance improvement over $G=5$, but it incurs a fourfold increase in per-step runtime, leading to very low sample efficiency in practice.
Low Sample Efficiency. GRPO computes sample advantage based on relative advantage. When the number of rollouts is small, it suffers from a severe issue of advantage vanishing, leading to a lack of effective training signals. However, as the number of samples increases, the relative advantage becomes diluted, making it harder for high-quality samples to exert sufficient influence. This results in low sample efficiency: although the computational overhead grows nearly linearly with the number of samples, the corresponding performance improvement is significantly diminished, as shown in Fig. 1.
Instability. During the later stages of training, the GRPO algorithm may exhibit signs of instability or collapse, characterized by a sharp decline in reward and sudden spikes in both gradient and policy entropy, as illustrated in Fig. 2. In extreme cases, this can escalate into a gradient explosion. Such instability presents a serious challenge to reliable long-horizon optimization, particularly in complex reasoning tasks where consistent gradient signals are critical for convergence. One contributing factor is the increased variance caused by the disappearance of relative advantages in the later stages of training. Supplementing with additional samples with non-zero advantages can help alleviate this issue.
((a))
((b))
((c))
Figure 2: When training Qwen2.5-VL-7B-Instruct on the Geo3k dataset using the GRPO algorithm, a collapse occurred.
4 Methodology
To address the issues in the previous section, we propose the EFRame, which consists of three main components: additional rollout, online filtering and replay buffer, as illustrated in Figure 3.
Figure 3: The overall workflow of EFRame builds upon the vanilla GRPO framework by introducing three key components: additional rollout, online filtering, and replay buffer.
4.1 Regular Rollout
For each batch of prompts, we begin with a regular rollout phase, following the original GRPO training process. This initial rollout serves two key purposes:
  • Obtain reward signals from regular (i.e., solvable) problems.
  • Identify more challenging prompts that fail to yield positive samples, which will be handled in subsequent stages.
Valuable samples are temporarily stored in the buffer for gradient updates.
4.2 Additional Rollout
GRPO computes loss over $G_{1}$ sampled outputs, making $G_{1}$ pivotal. Smaller $G_{1}$ limits learning capacity; while larger $G_{1}$ adds linear cost with limited benefit. For most questions, at least one positive sample can typically be found within a small number of rollouts. However, for more challenging problems, a larger number of samples and a higher temperature are often necessary to promote exploration and obtain meaningful reward signals.
To address this, we introduce an additional rollout mechanism: we perform an extra sampling round (e.g. $G_{2}$=20, temperature = 1.2) for challenging prompts.
Definition 4.1.
In GRPO-based algorithms with binary rewards (Premise, omitted in the following), difficult prompts are defined as those for which all sampled responses within the initial $G_{1}$ rollouts ($G_{1}$ is a smaller integer, such as 5) fail to obtain a positive reward, where $\bar{\mathcal{D}}\subseteq\mathcal{D}$ and:
$\bar{\mathcal{D}}=\left\{(q,a)~{}:~{}|\left\{\{o_{i}\}\sim\pi_{\theta_{old}(\cdot|q)}~{}:~{}\text{is\_equivalent}(a,o_{i})\right\}_{i=1}^{G_{1}}|=0\right\},$
The additional rollout is designed to activate deeper reasoning capabilities and increase the frequency of emergent “aha moments.”
4.3 Online filter
While additional rollout enhances the model’s ability to explore complex prompts, it also introduces computational overhead and potential training noise. In addition to filtering overlong samples, we design two additional filtering strategies to further enhance the stability and efficiency of the training process.
Regular Filter: In the first-round rollout, when the prompt is of regular difficulty, we adopt a conservative filtering strategy that retains all regular samples $O_{R}$ to balance efficiency and performance.
Definition 4.2.
Regular samples $O_{R}$ refer to any samples with a non-zero advantage on the prompt set $\mathcal{D}$:
$O_{R}=\left\{\{o_{i}\}\sim\pi_{\theta_{old}}(\cdot|q):(q,a)\sim\mathcal{D},\ \hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G_{1}})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G_{1}})}\neq 0\right\};$
At this stage, due to the moderate difficulty of prompts, all regular samples—both positive and negative—provide informative learning signals that contribute meaningfully to training.
Exploration Filter: When the prompt becomes challenging, we encourage the model to explore extensively in order to discover correct answers. Excluding samples with zero advantage, the remains can be categorized into the following two types.
Definition 4.3.
High-quality samples $O_{H}$ are positive samples for difficult prompts $\bar{\mathcal{D}}$:
$O_{H}=\left\{\{o_{i}\}\sim\pi_{\theta_{old}(\cdot|q)}:(q,a)\sim\bar{\mathcal{D}},\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G_{2}})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G_{2}})}>0\right\};$
low-quality samples are negative samples for difficult prompts $\bar{\mathcal{D}}$:
$O_{L}=\left\{\{o_{i}\}\sim\pi_{\theta_{old}(\cdot|q)}:(q,a)\sim\bar{\mathcal{D}},\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G_{2}})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G_{2}})}<0\right\}.$
Theorem 4.4.
For $o_{h}\in O_{H}$ and $o_{l}\in O_{L}$, $\sum\hat{A}_{h,0}+\sum\hat{A}_{l,0}=0$.
According to Definition 4.3 and Theorem 4.4, low-quality samples are significantly more numerous than high-quality ones, whereas the advantage values of high-quality samples are substantially higher. In such an exploration-driven setting, penalizing incorrect trajectories is not only inefficient but may also introduce noisy and unstable gradients. If gradients are computed from all samples, most of which are low quality, the informative signal from high-quality samples may be drowned out by chaotic updates from low-quality ones. To address this issue, we apply a more aggressive filtering strategy during the additional rollout—discarding all negative samples and retaining only positive samples for training. In practice, since this exploration phase targets only challenging prompts, the vast majority of generated samples (over 95%) are filtered out under this strategy.
To summarize, the policy gradient under our framework can be expressed as:
[…]
$\qquad\qquad\quad{}-\,\beta\,D_{\mathrm{KL}}\bigl{[}\pi_{\theta}\,\|\,\pi_{\mathrm{ref}}\bigr{]}\Bigr{)}\Biggr{]}.$
To prevent token-length bias, we compute the policy loss using the token-mean aggregation strategy.
4.4 Replay buffer
While GRPO is fundamentally an online RL algorithm, we find that experience replay with high-quality offline samples remains highly effective. In particular, trajectories discovered during the second-round rollout—are often of higher quality due to deeper exploration and carry rich reward signals. By selectively storing high-quality samples in the replay buffer, we amplify their influence during training, particularly when their gradients are frequently clipped due to low sampling probabilities. Furthermore, by enabling the model to repeatedly leverage these rare yet informative samples, experience replay plays a pivotal role in enhancing the stability of entropy.
Theorem 4.5 (Entropy Change under NPG Update).
If the actor policy $\pi_{\theta}$ is a tabular softmax policy updated via natural policy gradient (Kakade, 2001) with step size $\eta$. Then the change in policy entropy between two steps approximately satisfies:
${H}(\pi_{\theta}^{k+1})-\displaystyle{H}(\pi_{\theta}^{k})\approx-\eta\cdot\mathbb{E}_{s\sim d_{\mu}^{k}}\text{Cov}_{a\sim\pi^{k}_{\theta}(\cdot\mid s)}\!\bigl{[}\log\!\pi^{k}_{\theta}(a\mid s),\,A^{\pi^{k}}(s,a)\bigr{]},$
This theorem was first introduced by Liu (2025), and was organized and extended by Cui et al. (2025). Proof can be seen in Liu (2025) and Cui et al. (2025). ${H}$ indicates the policy entropy of policy model, and Cov denotes covariance, $\pi^{k}_{\theta}$ is the policy at step $k$, and $A^{\pi^{k}}(s,a)$ is the advantage function of action $a$ under state $s$. This result indicates that when high-advantage actions have low probability, they induce negative covariance, which increases entropy and reduces stability.
The high-quality samples identified through additional rollouts are typically responses to difficult prompts and thus are with relatively low probability. Using them only once yields limited and short-lived gradient effects, leading to an increase in model entropy. By incorporating these samples into the replay buffer, their probability of reoccurrence is effectively increased during training, enabling the model to transition from entropy increase to eventual convergence. In fact, the replay buffer here serves to reduce policy entropy; detailed discussions can be found in the following discussion section.
4.5 Discussion
The above subsections yield several practical insights regarding the interplay between Exploration-Filter-Replay:
Additional rollout: By prompting LLM to generate multiple responses to challenging queries ($x\in\bar{\mathcal{D}}$) under a high-temperature setting, the resulting positive samples are obtained with extremely low probability $\log\pi_{\theta}(y\mid x)$ and high advantage $\hat{A}(x,y)$. According to Theorem 4.5, this high-advantage, low-probability scenario rapidly increases the policy entropy $H$, thereby preventing the collapse of the entropy and maintaining effective exploration. At the same time, an excessive presence of low-quality samples can cause a nosedive in entropy, which is difficult to control. To address this, we filter out such samples and instead adopt a more controllable strategy to regulate entropy stability.
Replay Buffer: By storing high-quality positive samples within the replay buffer, the model can repeatedly utilize their gradient information for subsequent updates. As the model has already been updated on these (problem, answer) pairs $(x,y)$, the recalculated $\log\pi_{\theta}(y\mid x)$ becomes larger upon rehearing these pairs. According to the theorem, this shift leads to a reduction in policy entropy, preventing excessive growth in entropy.
5 Experiments
To validate the effectiveness of our methods, we demonstrate our experiment setup and the results in the following sections.
Table 1: Comparative results across multiple multimodal reasoning benchmarks indicate that our method achieves excellent performance.
Models Size MathVision MathVerse MathVista We-Math Average
Close-Source Models
GPT-4o Hurst et al. (2024) - 33.95 48.83 59.50 65.00 51.82
Claude-3.5-Sonnet Anthropic (2024) - 46.48 57.64 68.20 73.05 61.34
Open-Source Multi-Modal Large Reasoning Models
QvQ-Preview Bai et al. (2025) 72B 35.56 52.81 69.20 65.29 55.72
Skywork R1V Peng et al. (2025) 38B 39.31 40.91 67.50 60.06 51.95
Open-Source 7B Multi-Modal Models
MM-Eureka Meng et al. (2025) 7B 28.36 49.52 70.40 66.03 53.58
OpenVLThinker Deng et al. (2025) 7B 25.30 42.79 64.10 63.91 49.03
ThinkLite-VL Wang et al. (2025b) 7B 27.71 47.41 69.00 67.99 53.03
VLAA-Thinker Chen et al. (2025a) 7B 26.61 47.49 66.20 66.49 51.70
Qwen2.5-VL-7B-Instruct with RL
Qwen2.5-VL Bai et al. (2025) 7B 24.87 43.83 66.30 62.87 49.47
Qwen2.5-VL+GRPO 7B 29.11 47.51 72.60 67.53 54.19
Qwen2.5-VL+DAPO 7B 27.92 48.48 72.30 69.08 54.45
Ours 7B 28.82 (+3.95) 48.81 (+4.98) 73.40 (+7.10) 69.48 (+6.61) 55.13 (+5.66)
5.1 Experimental setup
Training Datasets. Our experiments are conducted on two datasets to evaluate the effectiveness of the algorithm in both in-domain and out-of-domain settings: (1) Geometry3K focused on geometric problems Lu et al. (2021a), (2) a 6k out-of-domain dataset randomly sampled from the dataset proposed in  Wei et al. (2025b; a), which randomly sampled from established open-source resources, including Geometry3K Lu et al. (2021a), GeoQA Chen et al. (2021), GeoQA-Plus Cao & Xiao (2022), Geos Seo et al. (2015), AI2D Kembhavi et al. (2016), TQA Kim et al. (2018), FigureQA Kahou et al. (2017), TabMWP Lu et al. (2022b), ChartQA Masry et al. (2022), IconQA Lu et al. (2021b), Clevr-Math Lindström & Abraham (2022), M3CoT Chen et al. (2024), and ScienceQA Lu et al. (2022a).
Evaluation Benchmarks. Our experiments consist of two parts. First, we evaluate the overall performance of the algorithms on four open multimodal math benchmarks: MathVision Wang et al. (2024a), MathVista Lu et al. (2023), MathVerse Zhang et al. (2024), and We-Math Qiao et al. (2024). These benchmarks cover diverse problem types such as geometry, charts, and tables, spanning multiple subjects and levels of visual reasoning difficulty. Second, we conduct a detailed analysis of algorithms on the Geometry3K dataset to examine the algorithmic behaviors.
Baselines. We compare our models with a range of advanced multimodal large language models. These include GPT-4o Hurst et al. (2024) and Claude-3.5-Sonnet Anthropic (2024), which are state-of-the-art closed-source MLLMs. QvQ-Preview-72B[2] and Skywork R1V-32B  Peng et al. (2025) are multimodal large reasoning models specifically designed for R1-like chain-of-thought capabilities. At the 7B scale, we compare with MM-Eureka-7B Meng et al. (2025), OpenVLThinker-7B Deng et al. (2025), ThinkLite-VL-7B Wang et al. (2025b), and VLAA-Thinker-Qwen2.5VL-7B Chen et al. (2025a). All of these 7B models are based on Qwen2.5-VL and are enhanced for mathematical reasoning via fine-tuning, reinforcement learning, or iterative self-improvement techniques. In addition, we compare the performance of GRPO and DAPO on the Qwen2.5-VL-7B-Instruct model to assess the impact of different strategies under consistent model and training settings.
Implementation details. Our work is based on EasyR1 and VeRL frameworks Yaowei et al. (2025); Sheng et al. (2025). We initialize all policy models with Qwen2.5-VL-7B-Instruct and conduct experiments on 8 A800 GPUs (40GB). Unless otherwise specified, we follow the default EasyR1 settings: maximum response length of 2048, global batch size 128, rollout batch size 512, rollout $G_{1}$=5, temperature 1.0, and learning rate $10^{-6}$. GRPO and our method adopt $\epsilon_{low}=\epsilon_{high}=0.2$ and a binary reward to avoid relative advantage bias (see detail in Appendix A), while DAPO includes $\epsilon_{high}=0.3$ and an additional length-aware linear penalty (up to 0.5) for responses exceeding 1536 tokens. For our method, we perform additional rollouts with $G_{2}=20$, temperature = 1.2; responses with absolute advantage larger than 1 are stored in a replay buffer of size 5120, with experience replay triggered every 5 steps.
Table 2: The performance comparison on the Geometry3K dataset shows that our method achieves a significant improvement over existing approaches.
Models Qwen2.5-VL-Instruct Ours GRPO DAPO
Accuracy 38.44 55.24(+16.80) 50.75 (+12.31) 49.09 (+10.65)
Time/step(s) - 442($\times$1.35) 327($\times$1.00) 182($\times$0.57)
5.2 Main results
((a))
((b))
((c))
Figure 4: On the Geometry3K dataset, our method not only achieves excellent performance but also demonstrates superior exploration capability and training stability.
Performance. In our experiments, we employ the Exploration-Filter-Replay framework, yielding notable improvements on both out-of-domain and in-domain datasets. As shown in Table 1, the proposed method delivers consistently strong results across four multimodal reasoning benchmarks, with an average gain of +5.66 over the base model (Qwen2.5-VL-7B-Instruct), outperforming GRPO by 19.9%. On the in-domain Geometry3K dataset, the improvement is even more pronounced: Table 2 reports a +16.80 gain, representing a 36.5% relative advantage over GRPO.
Exploration. As shown in Fig. 4, our framework maintains the policy entropy around 1, striking a balance between sufficient exploration and exploitation. In the later training stages, when the vanilla GRPO tends to suffer from vanishing advantages, our method further increases entropy to encourage continued exploration of informative reward signals.
Stability. Our framework leverages additional rollouts to enrich reward signals, while an online filtering mechanism mitigates training variance by excluding low-quality samples. Extensive experiments reveal that the proposed method circumvents collapse issues—occasionally encountered in GRPO and DAPO. Furthermore, the steadily increasing response length during training reflects the model’s progressively improved reasoning ability.
5.3 Entropy Control
To achieve a well-balanced exploration process, we incorporate entropy control into EFRame by adjusting two key hyperparameters: the additional rollout temperature $t_{a}$ and the replay buffer size $R_{s}$. Specifically, increasing the temperature in the additional rollout phase encourages broader exploration, enabling the model to discover low-probability but potentially high-reward trajectories. At the same time, we increase the $R_{s}$ to promote convergence and prevent excessive entropy growth. Together, these mechanisms allow for a controlled yet sufficiently rich exploration process—ensuring that the model explores meaningfully without destabilizing training due to overly random behavior.
((a))
((b))
((c))
((d))
((e))
((f))
Accuracy $t_{a}=1.0$ $t_{a}=1.1$ $t_{a}=1.2$
$R_{s}=5120$ 53.74 52.41 55.24
$R_{s}=2560$ 52.75 54.41 52.75
$R_{s}=0$ 52.41 51.08 50.25
Entropy $t_{a}=1.0$ $t_{a}=1.1$ $t_{a}=1.2$
$R_{s}=5120$ 0.089 0.305 0.726
$R_{s}=2560$ 0.091 0.312 2.366
$R_{s}=0$ 0.092 2.437 4.288
Figure 5: We report the peak accuracy within the first 200 training steps and the average entropy over the first 150 steps. Experiments on the GEO3K dataset using Qwen2.5-VL-7B-Instruct under different settings of additional rollout temperature $t_{a}$ and replay buffer size $R_{s}$ demonstrate that increasing $t_{a}$ enhances exploration, reflected in higher entropy, while enlarging $R_{s}$ promotes convergence, leading to lower entropy. Optimal performance is achieved when exploration and convergence are well-balanced.
We observe in Fig. 5 that increasing the additional rollout temperature encourages broader exploration, leading to a rise in policy entropy. In contrast, enlarging the replay buffer size, which increases the number of high-quality samples reused during training, promotes convergence and results in decreased entropy. These two hyperparameters exhibit opposing effects on entropy dynamics, providing a practical mechanism for entropy control within EFRame. We find that model performance peaks when the entropy stabilizes around a moderate level. Fig. 5 reports the accuracy under different $t_{a}$ and $R_{s}$, highlighting the impact of entropy control of EFRame on learning efficacy. Our experiments demonstrate that entropy control can be achieved with high precision by combining high-temperature positive sampling and experience replay, which are more effective than using entropy-based reward functions.
6 Ablation Studies
To assess the contribution of each component in our framework, we conduct an ablation study to examine whether all three modules—additional rollout, online filtering, and replay buffer—are essential for achieving the observed improvements in stability and overall performance.
6.1 Additional Rollout
Disabling the additional rollout module exposes the model to a fundamental trade-off inherent in GRPO-based training. When the number of samples $G_{1}$ is small, the reward signal is constrained by the base model’s pass@$G_{1}$, limiting the model’s ability to solve challenging problems and suppressing the emergence of “aha moments.”
Claim 1.
High-quality samples have the potential to activate deeper reasoning abilities within the model.
As shown in Fig. 6, the model’s reward on the training set plateaus at a relatively low level, restricting further gains in test accuracy. In the early stages of training, the entropy remains low, indicating limited exploration. However, once the model hits this performance bottleneck, entropy begins to rise—yet this delayed exploration yields minimal returns, suggesting that the model struggles to escape local optima. The additional rollout module is specifically designed to break this situation, enabling targeted deep exploration.
((a))
((b))
((c))
Figure 6: The ablation study conducted on the Geometry3K dataset not only validates the effectiveness of our proposed method but also reveals the distinct roles played by samples of varying quality during the training process.
6.2 Online Filtering
Removing the online filtering module results in significant instability and inefficiency during training. Without filtering, the model treats all samples—regardless of their advantage—as equally important, allowing zero-advantage samples to increase gradient variance and low-quality negative trajectories to distort the learning signal. This problem is further exacerbated during the additional rollout phase, where abundant incorrect samples significantly amplify gradient variance, making it difficult for the model to retain meaningful patterns from rare high-quality responses. As a result, training becomes slower, more chaotic, and more prone to divergence.
Claim 2.
Low-quality samples from challenging prompts often mislead optimization, causing the model to converge prematurely to suboptimal solutions and impeding further learning.
As shown in Fig. 6, during the early stages of training, the inclusion of a large number of negative samples significantly boosts the model’s performance on the training set but yields little improvement on the test set. This indicates that most low-quality samples lead the model to converge more quickly to a local optimum, offering little benefit to its generalization ability. Moreover, the forward and backward computation of these negative samples incurs substantial overhead, reducing the sample efficiency of the overall training process. Online filtering plays a crucial role in maintaining gradient clarity and training stability by selectively preserving informative samples and discarding neutral or low-impact ones.
6.3 Replay Buffer
When the replay buffer is disabled, we observe a drastic drop in performance, even falling below that of the vanilla GRPO baseline, as shown in Fig. 6. This counterintuitive result reveals that the abundant high-quality samples generated during the additional rollout phase can actually become detrimental without a proper replay mechanism.
Claim 3.
While high-quality samples may induce high-entropy exploration, incorporating experience replay enables the model to retain and consolidate useful gradients, thus restoring training stability.
As high-quality samples typically have low generation probabilities, rewarding them shifts the model’s response distribution toward low probability but correct outputs, disrupting the convergence trajectory and consequently increasing entropy. Due to their high advantage values and low sampling probabilities, high-quality samples are more likely to have their gradients clipped during training. In the absence of a replay buffer, such valuable samples contribute only short-lived and unstable gradients that are quickly overwritten. Instead of reinforcing beneficial behavior, this transient gradient signal increases policy uncertainty, pushing the model into a state of chaotic, high-entropy exploration, ultimately harming both learning stability and performance. To address this issue, we leverage experience replay to ensure that high-quality samples continuously provide effective gradients, enabling the model to converge along the new trajectory.
7 Conclusion and Discussion
In this work, we investigate the impact of sample quality on RL for large language models, with a particular focus on addressing challenging prompts. Our findings show that low-quality incorrect samples—those with low absolute advantage—tend to trap the model in local optima and offer little or even negative contribution to training. In contrast, high-quality samples with large advantages play a critical role in guiding the model through a meaningful transition from entropy-driven exploration to entropy-reducing convergence. This process is essential for eliciting deeper reasoning capabilities within the model. To this end, we propose an Exploration-Filter-Replay framework that emphasizes the discovery, selection, and sustained utilization of high-quality samples. Our method introduces additional rollouts to enhance reward density, employs online filtering to suppress noisy samples, and integrates experience replay to amplify the impact of high-quality samples across training iterations. This strategy not only helps transform isolated “aha moments” into a sustained “aha state” but also shifts reinforcement learning from a “reinforcement lottery” toward a more structured and reliable process of reinforcement learning.
Appendix
Appendix A Relative Advantage Bias of GRPO
In GRPO, computing relative advantages within a group can introduce bias, especially when the reward becomes multi-dimensional. For example, as shown in Fig. 7, if the reward comprises both accuracy and auxiliary signals such as length, format, or entropy, a significant issue arises when the accuracy signal vanishes within a group—i.e. when all responses are either correct or incorrect. In such cases, auxiliary rewards (e.g., length) dominate the advantage computation. This may result in incorrect responses being rewarded simply for being shorter, or correct responses being penalized for being longer. Such misleading gradients introduce irreversible noise into training and lead to reward hacking. Therefore, in the training involved in this paper, we use binary rewards whenever possible.
Figure 7: During training with GRPO, when the primary reward signal (accuracy) vanishes, other auxiliary rewards become dominant, which introduces bias.